[Brahms-l] Shift report 20050419 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Apr 19 2005 - 08:02:59 EDT
Shifter leader was SB
Shifters were 


1. Store from previous shift continued until 3:00 am. New store from 
   5:00 am. 4.25 hours of physics running

2. Runs 14116,14117 with FS @1723A
   Runs 14118,14119,14120,14122,14123 With FS @861A

3. TPM2 was left out of run during previous shift due to
    unknown problems. The problem continued into the night shift
   so TPM2 was out the whole night.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Apr 19 08:04:01 2005

This archive was generated by hypermail 2.1.8 : Tue Apr 19 2005 - 08:04:19 EDT